# Sim-DETR: Unlock DETR for Temporal Sentence Grounding

### v1

@@ 5.2 Comparison with State-of-the-Art Methods
Results on QVHighlights are shown in Table 1. Compared to the latest published work, BAM-DETR [25], our Sim-DETR outperforms it by an average of +2.88% across all metrics. More importantly, it achieves remarkable gains of +4.93% and +2.27% in R1@0.5 and R1@0.7, respectively. In comparison TR-DETR [57], which uses standard DETR [2] as its basic framework and is used as our baseline model, our Sim-DETR achieves an average improvement of +4.31% and +4.41% in mAP for test and validation, respectively. These improvements underscore the effectiveness of our Sim-DETR in fundamentally addressing conflicts within and between queries. We further conduct a comparison with SpikeMba [31], a contemporary method built on the Mamba [17] framework. Without any complex module designs, the consistent improvements across all metrics highlight the simplicity and effectiveness of our Sim-DETR. Sim-DETR establishes an efficient baseline for the task, significantly advancing research in TSG.
Results on Charades-STA and TACoS are presented in Table 2, where we consistently outperform all methods on both datasets. Compared to the state-of-the-art methods, we achieve a +1.97% improvement on the Charades-STA dataset (with SpikeMba as the SOTA) and a +3.89% improvement on the TACoS dataset (with CG-DETR [46] as the SOTA). These enhancements further demonstrate the outstanding capability of our method in temporal localization. More importantly, they confirm that our Sim-DETR exhibits significant performance advantages across various scenarios, types, and lengths of videos, highlighting its robustness and generalizability.
@@
Results with Different Backbones. Following previous works, we conduct two additional experiments on the QVHighlights dataset. These experiments use VGG and I3D for visual feature extraction and use Glove [48]embedding as text features. Table 3 presents the detailed results of these two experiments, where we achieve state-of-the-art results in both cases. Notably, for the video feature extractor I3D, we surpass the latest published method by +4.26%. These results further demonstrate the excellent generalization capabilities of our Sim-DETR across different backbones.
@@ 5.3 Ablation Study
Roadmap for building a simple yet effective TSG baseline is presented in Table 5. We employ TR-DETR without MR2HD module as our baseline, which achieves an mAP of 44.97%. Since our Query Grouping and Ranking (QGR) utilizes $\mathcal{L}_{\text{iou}}$ proposed by [25, 58], we additionally conduct an ablation study to isolate the impact of $\mathcal{L}_{\text{iou}}$. (1) QGR: To mitigate conflicts and optimization challenges arising from overly similar semantics between queries, we introduce the QGR, which improves the baseline by +2.07% on mAP, representing a subtle yet significant enhancement. (2) Global-Local Bridging (GLB): To address conflicts and ambiguities between global semantics and local localization within queries, we introduce GLB to enable seamless transformation through query-to-frame alignment. Experimental results show a performance boost to 48.17% on mAP, indicating the effectiveness of GLB in bridging global-to-local transformations. However, without QGR to address semantic conflicts in temporal DETR, the recall predictably decreased by -3.1%. (3) Overall Framework: With these adjustments, our Sim-DETR achieves a remarkable SOTA performance: 49.50% on mAP and 69.48% on R1@0.5.

### v2

@@ 5.2 Comparison with State-of-the-Art Methods
Results on QVHighlights are shown in Table 1. Compared to the latest published work, BAM-DETR [lee2025bamdetr], our Sim-DETR outperforms it by an average of +2.88% across all metrics. More importantly, it achieves remarkable gains of +4.93% and +2.27% in R1@0.5 and R1@0.7, respectively. In comparison TR-DETR [sun2024trdetr], which uses standard DETR [carion2020detr] as its basic framework and is used as our baseline model, our Sim-DETR achieves an average improvement of +4.31% and +4.41% in mAP for test and validation, respectively. These improvements underscore the effectiveness of our Sim-DETR in fundamentally addressing conflicts within and between queries. We further conduct a comparison with SpikeMba [li2024spikemba], a contemporary method built on the Mamba [gu2023mamba] framework. Without any complex module designs, the consistent improvements across all metrics highlight the simplicity and effectiveness of our Sim-DETR. Sim-DETR establishes an efficient baseline for the task, significantly advancing research in TSG.
Results on Charades-STA and TACoS are presented in Table 2, where we consistently outperform all methods on both datasets. Compared to the state-of-the-art methods, we achieve a +1.97% improvement on the Charades-STA dataset (with SpikeMba as the SOTA) and a +3.89% improvement on the TACoS dataset (with CG-DETR [moon2023cgdetr] as the SOTA). These enhancements further demonstrate the outstanding capability of our method in temporal localization. More importantly, they confirm that our Sim-DETR exhibits significant performance advantages across various scenarios, types, and lengths of videos, highlighting its robustness and generalizability.
@@
Results with Different Backbones. Following previous works, we conduct two additional experiments on the QVHighlights dataset. These experiments use VGG and I3D for visual feature extraction and use Glove [pennington2014glove]embedding as text features. Table 3 presents the detailed results of these two experiments, where we achieve state-of-the-art results in both cases. Notably, for the video feature extractor I3D, we surpass the latest published method by +4.26%. These results further demonstrate the excellent generalization capabilities of our Sim-DETR across different backbones.
@@ 5.3 Ablation Study
Roadmap for building a simple yet effective TSG baseline is presented in Table 5. We employ TR-DETR without MR2HD module as our baseline, which achieves an mAP of 44.97%. Since our Query Grouping and Ranking (QGR) utilizes $\mathcal{L}_{\text{iou}}$ proposed by [lee2025bamdetr, sun2024rgtr], we additionally conduct an ablation study to isolate the impact of $\mathcal{L}_{\text{iou}}$. (1) QGR: To mitigate conflicts and optimization challenges arising from overly similar semantics between queries, we introduce the QGR, which improves the baseline by +2.07% on mAP, representing a subtle yet significant enhancement. (2) Global-Local Bridging (GLB): To address conflicts and ambiguities between global semantics and local localization within queries, we introduce GLB to enable seamless transformation through query-to-frame alignment. Experimental results show a performance boost to 48.17% on mAP, indicating the effectiveness of GLB in bridging global-to-local transformations. However, without QGR to address semantic conflicts in temporal DETR, the recall predictably decreased by -3.1%. (3) Overall Framework: With these adjustments, our Sim-DETR achieves a remarkable SOTA performance: 49.50% on mAP and 69.48% on R1@0.5.